10L of hard water gives 0.56g of lime for removing hardness .Hence temporary hardness which is due to Ca(HCO3)2 in ppm of CaCO3 is:

Grade:12

1 Answers

Pankaj
9 years ago
Total volume of water = 10L
Desnity of water = 1 Kg/L
Mass of 10 L water = 10kg
Molar mass of water = 10 kg
Number of moles of water = 10000/18 = 555.55
Molar mass of Ca(HCO3)2 = 162 g
Number of moles ofCa(HCO3)2 = 0.56/162 = 0.003456

ppm = \frac{3.45\times 10^-^3}{555.55}\times 10^6 =6.22